# GraphEx: A Graph-based Extraction Method for Advertiser Keyphrase Recommendation

### v2

@@
Abstract.
Online sellers and advertisers are recommended keyphrases for their listed products, which they bid on to enhance their sales. One popular paradigm that generates such recommendations is Extreme Multi-Label Classification (XMC), which involves tagging/mapping keyphrases to items. We outline the limitations of using traditional item-query based tagging or mapping techniques for keyphrase recommendations on E-Commerce platforms. We introduce GraphEx, an innovative graph-based approach that recommends keyphrases to sellers using extraction of token permutations from item titles. Additionally, we demonstrate that relying on traditional metrics such as precision/recall can be misleading in practical applications, thereby necessitating a combination of metrics to evaluate performance in real-world scenarios. These metrics are designed to assess the relevance of keyphrases to items and the potential for buyer outreach. GraphEx outperforms production models at eBay, achieving the objectives mentioned above. It supports near real-time inferencing in resource-constrained production environments and scales effectively for billions of items.
@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systemsccs: Information systems Data miningccs: Information systems Sponsored search advertisingccs: Mathematics of computing Graph algorithmsccs: Computing methodologies Information extraction
1. Introduction
In the online e-commerce advertisement space, keyphrase recommendation is offered to sellers/advertisers who want to bid on buyers/users’ search queries for a better placement of their inventory on the search result pages (SRP) which increases their engagement. Keyphrases are generally recommended as shown in figure 1 in real time for the items if they are relevant to them. Advertisers only want to bid on keyphrases that are actual queries and not queries that seem plausible but non-existent for targeting purposes. Since the nature of the problem is mapping items to multiple queries to increase the potential reach of advertisers, this problem of keyphrase recommendation can be formulated as an Extreme Multi-Label Classification (XMC) problem, see  (Dahiya et al., 2023a; Jain et al., 2023; Dahiya et al., 2023b; You et al., 2019; Dahiya et al., 2021; Mishra et al., 2024). The data for training the XMC models is sourced from search logs that document the items shown to buyers when they input search queries [^1] We term buyer search queries as keywords. And use keywords and keyphrases interchangeably.. A keyword and an item are paired together and become a part of the training dataset when they co-occur in the search logs a certain number of times with significant buyer interactions on those items. XMC models in the context of keyword recommendation are supervised tagging techniques that map items to keyphrases.
@@ 1. Introduction
Figure 1. Screenshot of our keywords for manual targeting in Promoted Listings Priority for eBay Advertising.
The impact of XMC tagging models can be determined by how often each keyphrase is searched. Keyphrases can be classified as head or tail keyphrases according to their search frequency. Head keyphrases are generally less in number but searched frequently by buyers. Targeting such head keyphrases leads to increased revenue since more buyers inclined to search for them, resulting in more clicks and more buys. XMC models focus on recommending more tail keyphrases (Dahiya et al., 2023b, a) and avoid the head keyphrases despite them being relevant to the item. In addition, tail keyphrases/queries will not have too many items relevant to it, so the chances of the relevant seed item getting proper positioning through organic retrieval without any promotion increases — making them unappetizing to advertisers.
There are various caveats to recommending keyphrases using XMC tagging formulation trained on engagement data from search logs. Search logs with engagement data is highly skewed as shown in Figure 2, with 90% of items associcated with only one query in terms of clicks/sales. This absence of engagement (clicks/sales) doesn’t necessarily mean that the keyphrase/query is irrelevant to the item. When buyers are presented items in relation to the query/keyphrase, the presentation is biased as the items are ranked and their ranking can affect buyer engagement. This biased presentation implies that just because an item doesn’t have clicks/sales in relation to the keyphrase doesn’t mean that the item is irrelevant to the keyphrase. It could mean that the item was not popular and was retrieved at a lower rank for the query in question and influenced buyers to ignore it. These unpopular items need the help of advertisement to level the playing field by promoting their items to a favourable rank to garner more buyer engagement. Thus traditional XMC models which rely on modelling item-query interaction inherit this bias towards popular items which are ranked at the top and miss the targeting for non-popular items which constitute 90% of the inventory and are the main focus of advertisement.
While this bias has been discussed in  (Joachims et al., 2007; Yue et al., 2010; Joachims et al., 2017; Deffayet et al., 2023), we contextualize it in this domain of advertiser keyphrase recommendation. Continuing from the limitations of the data on which XMC tagging models are trained, the same training set and bias perpetuate the lack of diversity of different XMC models. Even with a $10\%$ increase in the precision/recall scores of subsequent XMC models, the recommendations do not have sufficient diversity to obtain substantial clicks. In addition the lack of ground-truth for most items leads to inaccurate offline evaluation which is conventionally ground truth-based (Precision/Recall). We offer a diverse set of alternate evaluation metrics which are comprehensive in measuring a model’s performance and diversity for these skewed distributions and test it in an online setting.
Figure 2. Distribution of data in ecommerce.
1.1. Scope and Contributions
@@
In this work, we limit ourselves to retrieving keyphrases based on item’s title and the keywords from the items’s categorical populace, especially those keyphrases that are actively and frequently searched by buyers. The extraction is done in an unsupervised setting where the keyphrases for each item are unknown during training. In fact, we restrict the curation of keyphrases (more details in 3.2) to include only those that have high search volume (number of searches made by buyers) based on buyer searches. XMC models suffer from item popularity biases, which can result in a severely restricted set of keyphrases for non-popular items [^2] Non-popular items constitutes more than 80% of total items or recommending a lot of tail keyphrases for items. Our distinction is that by decoupling the keyphrases from item engagement and using categorical population dynamics of keyphrases, we keep the essential bias towards head keyphrases (attractive to advertisers) while getting rid of the negative bias (against non-popular items which are the main target of advertisement).
The models are required to be frequently refreshed, i.e., the models are regularly trained and tuned on recent data to keep up with the latest buyer search queries. Thus, models with smaller training and setup times are necessary, and minimal tuning is crucial to decrease the engineering effort. In addition, due to the ever changing buyer query space, frequent model refreshes is required accommodate newer keywords. For e-commerce platforms, it is also vital that the recommendations are in real-time or near real-time, so the models should also have inference latency of a few milliseconds.
In this work, our aim is to overcome the many challenges as described above. Our contribution to this work is summarized as:
@@
  • An innovative graph-based extraction algorithm for keyphrase recommendation that is simple, transparent and easy to interpret.
@@
  • The design of the algorithm and the process of data collection have been specifically geared towards mitigating item popularity bias while maintaining advertiser friendly head keyphrase bias.
@@
  • Provide a new robust framework for evaluation of incremental impact of recommendation models in terms of performance and diversity metrics.
  • A scalable and sustainable model that runs without GPUs for hundreds of millions of items in real-time.
2. Related Work
@@
Keyphrase generation via open-vocabulary models like GROOV (Simig et al., 2022), One2Seq (Yuan et al., 2020; Xie et al., 2022) and One2One (Meng et al., 2017; Chen et al., 2019a, b) are susceptible to recommending keyphrases that are not part of the label space. Another formulation for keyphrase recommendation is keyphrase extraction with methods such as keyBERT(Grootendorst, 2020), which have conventionally treated keyphrase recommendation as a two-step problem: keyphrase generation and keyphrase ranking. The basic keyBERT module considers keyphrase generation as an n-gram based permutation problem, i.e., it generates all possible n-grams for a given n-gram range. The keyphrase ranking module then orders them using an encoder-based ranker tuned on some domain-specific supervised signal. This simple generation framework presents two main issues: 1) the token space is limited by token adjacency and token presence in item’s text 2) the keyphrase should also be in the universe of queries that buyers are searching for; which this simple generation model does not ensure.333keyBERT can also use LLMs as generators, but their time complexity is substantial. We limit our comparison to models with shorter training and inference times based on the scope discussed in section 1.1. The models deployed at eBay fulfill these requirements and provide good recommendations. We choose 4 representative models from the deployed ones; fastText, Graphite, Rules Engine (RE) and SimilarListing(SL) variants.
fastText (Bojanowski et al., 2017; Joulin et al., 2017) is a basic linear neural network model that generates word vectors with CBOW architecture and uses a single classification layer to generate predictions. It uses hierarchical softmax for faster training on a large number of classes (keyphrases) and subword embeddings for better representation and inferencing. Graphite (Mishra et al., 2024) is the state-of-the-art fastest XMC model that uses bipartite graphs to map words/tokens to the data points and then map them to the labels associated with the data points. It is implemented for multi-core systems having infinitesimal training time and uses parallelization for real-time inferencing. Both fastText and Graphite are deployed at eBay. There are four other proprietary models also deployed at eBay for keyphrase recommendation. Rules Engine (RE) is a simple technique that stores item-keyphrase associations based on their cooccurrences (associated with buyer activity) in the search logs during the last 30 days. It recommends keyphrases only for items in which buyers have shown interest and not for any new items. Likewise, RE-trank recommends only the queries that were ranked in the top slots for the existing association of items. We also compare with a few other variants that are based on similar listings (SL) and their related queries. These are; SL-query that determines the existing listings/items which share some keywords, thus recommending other’s associated keywords, and lastly SL-emb (Wang et al., 2021) which uses embeddings of the item’s title to compare and find similar listings then recommend the related queries. SL-emb is a dense retrieval model whose inference is implemented in two stages, embedding generation and ANN(Malkov and Yashunin, 2020). Note that of the SL models only SL-emb can handle cold start conditions on new items such as fastText and graphite. The implementation of the RE and SL techniques also employs a few other methods which we can’t discuss due to proprietary constraints.
3. GraphEx Model
@@ 3.1. Problem Formulation
For efficiently solving the recommendation problem, we use the formulation of a permutation problem that permutes the title strings to match a given set of keyphrases. Let’s consider a title string with $l$ words in it. The goal is to generate permutations of differing lengths from the $l$ words. Now, given a list of predefined keyphrases, the possible permutations of the title string is constrained to match the keyphrases. Therefore, each permutation can exactly match a keyphrase or be part of some keyphrases, but if a title token isn’t part of any keyphrase then it is ignored. Thus, it does not limit the permutations to token adjacency or token presence in item’s text. A naive brute force method is to generate all possible permutations of the $l$ words which will take $O(l!)$ time. Each keyphrase can be validated using hashing and string comparisons (each word can be an integer) and thus can take overall $O(l\times l!)$ time. This is infeasible to perform in real-time with limited amount of resources.
3.2. Dataset Curation
@@
We aggregate our datasets from the search logs generated during buyer sessions on eBay.com. The keyphrases that buyers input during the search sessions are curated based on certain criteria, which we discuss here. eBay’s search engine Cassini shows a sufficient number of items (Recall Count) for each input query in its search results. Cassini determines the Leaf Category of the keyphrase and it is the same as the top-ranked item’s leaf category (lowest-level product categorization). We restrict the number of curated keyphrases by only considering those that are heavily searched by the buyers. The number of times a keyphrase is queried is termed as (Search Count). Due to proprietary reasons, we cannot further delve into the details. The absolute values of Recall Count and Search Count are not essential in fact, an anonymized ranking works well. All the unique keyphrases are aggregated for each Top level category (metacategory) and are grouped for each Leaf Category within the metacategory. Each keyphrase is associated with a Search Count and a Recall Count. Note that a keyphrase can be duplicated across different Leaf Categories.
3.3. Terms and Notations

### v3

@@
Abstract
Online sellers and advertisers are recommended keyphrases for their listed products, which they bid on to enhance their sales. One popular paradigm that generates such recommendations is Extreme Multi-Label Classification (XMC), which involves tagging/mapping keyphrases to items. We outline the limitations of training XMC models on click data for keyphrase recommendations on E-Commerce platforms. We introduce GraphEx, an innovative graph-based approach that recommends keyphrases to sellers using extraction of token permutations from item titles. Additionally, we demonstrate traditional metrics such as precision/recall isn’t reliable on click-based data in practical applications, thereby necessitating a robust framework to evaluate performance in real-world scenarios. Our evaluation is designed to assess the relevance of keyphrases to items and the potential for buyer outreach. GraphEx outperforms production models at eBay, achieving the objectives mentioned above. It supports near real-time inferencing in resource-constrained production environments and scales effectively for billions of items.
Index Terms:
@@ Abstract
I Introduction
In the online e-commerce advertisement space, keyphrase recommendations are offered to sellers/advertisers who want to bid on buyers/users’ search queries for a better placement of their inventory on the search result pages (SRP) which increases the item’s engagement. Keyphrases [^1] We term buyer search queries as keyphrases and use queries and keyphrases interchangeably. are generally recommended as shown in Figure 1(b) on the right in real time for the items if they are relevant to them. The keyphrase recommendations provided by Advertising are then matched to actual queries by eBay Search and enter auctions where the sellers/advertisers with the highest bid wins the auction and gain a prominent sponsored placement on the SRP page. Hence, it is important to suggest keyphrases that are an exact match to the queries to prevent missed targeting. When buyers interact with the SRP by searching for a query and clicking on an item as shown in Figure 1(a), the advertiser or seller is charged (CPC or cost-per-click business model) for their advertised item. Any interaction (click, add to cart, buy etc.) on the SRP page is logged in the search logs. An association between the search queries or keyphrases and items can be derived which we call as click-based ground truths.
(a) Buyer side
(b) Seller Side
@@ I Introduction
Figure 1: Screenshot of our keyphrases for manual targeting in Promoted Listings PriorityTM for eBay Advertising.
This problem of keyphrase recommendation has been formulated as an Extreme Multi-Label Classification (XMC) problem and well studied in  [1, 2, 3, 4, 5, 6]. The data for training the XMC tagging models is generally the click-based ground truths sourced from search logs. There are multiple challenges impacting this research area which we describe in detail.
I-A Challenges
I-A1 Budgeted Recommendation
Modern recommendation systems operate on a budget of recommendations, i.e. a maximum number of recommendations (which in eBay’s case is 1000 keyphrases per adgroup with an adgroup containing at max 1000 items). Within this constrained budget@k [7] the targeting significance of every keyphrase becomes paramount. Keyphrases can be classified as head or tail keyphrases according to their search frequency. Head keyphrases are generally less in number but searched frequently by buyers. Targeting such head keyphrases leads to increased revenue since more buyers are inclined to search for them, resulting in more clicks and more buys. XMC models are agnostic towards head or tail keyphrases and end up focusing on recommending the more copious tail keyphrases [3, 1] while on a budget — missing out on potentially more important head keyphrases.
I-A2 Click data biases
The large set of labels or keyphrases in the click data challenges exhaustive annotation, resulting in missing labels due to sparsity of dataset (i.e. 96% of items don’t have clicks associated with them) and the clicks of items are influenced by various biases like popularity bias, exposure bias [8], sample-selection bias [9] etc. While clicks can be treated as reliable labels for relevance, the absence of clicks cannot be taken as a sign of irrelevance [10].
While these biases have been discussed in  [11, 12, 13, 14], we contextualize them in this domain of advertiser keyphrase recommendation. When buyers are shown items based on a specific query or keyphrase, the way the presented items are ranked can introduce bias, influencing buyer engagement. This biased ranking suggests that an item lacking clicks or sales for a given keyphrase isn’t necessarily irrelevant to that keyphrase [15]. Instead, it might be less popular and thus lower ranked, leading buyers to overlook it (exposure and popularity biases). These unpopular items need the help of advertising to level the playing field by promoting their items to a favorable rank, increasing their visibility and engagement. In explicit feedback data, such as click-based data, signals can be Missing-not-at-Random (MNAR) [10], which means that systems trained solely on these feedback signals are likely to perpetuate these biases in their predictions.
Even for the items that get a click, 90% of such items are associated with only one query in terms of clicks — as shown in Figure 2, while sellers expect a healthy set of recommendations from us (around 20-30 per item).
Figure 2: Distribution of click-data in terms of items and the number of queries associated with them.
I-A3 Click-based Evaluation
Sourcing the same click data as the training set for different XMC tagging models perpetuates not only the described biases but also the lack of diversity in the model’s recommendations. The offline evaluation of such models is typically done using metrics like Precision, Recall, F1 and so on. These metrics facilitate comparison by emphasizing retrieval capability, i.e how well are the models able to retrieve the existing associations of keyphrases with the items.
The diversity issue is exacerbated by evaluating with these metrics due to two main reasons: Lack of Ground Truths and Model Convergence. Due to the curation process, there is a lack of ground truths, since most items are excluded from the training set, as evident from Section I-A2 and Figure 2. So, the metrics aren’t able to evaluate a model’s recommendation beyond the labels gathered per item which are very sparse. This is problematic especially considering the budgetary constraints described in Section I-A1.
To understand the convergence of the model, let us take an example instance $T$, associated with a ground truth label $k1$ in the training set. All candidate models make certain choices to increase the probability of predicting $k1$ for inputs similar to $T$. This aligns the tagging-based models to predict a similar subset of labels for $T$, thus reducing diversity between the predictions of different models. Even with a $10\%$ increase in the precision/recall scores of subsequent XMC models, the recommendations do not have sufficient diversity to obtain substantial clicks. The impact of models is further dampened by the 100% recall models in production (database lookup of queries that generated clicks in relation to the items). Models with higher recall will have less impact as they will be de-duplicated against the 100% recall model’s recommendations.
I-A4 Keyphrase Targeting
The XMC tagging models are required to be regularly updated (preferably daily) to keep up with the churn of new queries (2% churn every day) and other factors such as holidays and seasonality. Out of Vocabulary (OOV) models can recommend keyphrases that are absent in the training set — thus circumventing the need for daily training.
The OOV models however also suffer from biases in click data as described in [7]. Data augmentation techniques such as rejection sampling [7] have been shown to mitigate these biases. As described earlier, the exact matching of recommended keyphrases to the search queries in auctions along with the budgetary constrains from section I-A1 makes OOV model’s inaccurate targeting less desirable.
Disengaging from the click-based associations can help mitigate their biases, however the targeting and the evaluation of such models become a challenge.
I-A5 Execution Performance
For e-commerce platforms, a vital necessity is that the recommendations are in real-time or near real-time, so the models should also have inference latency of a few milliseconds. Alternatively, models should have a sufficiently low inference latency for daily batch prediction. Due to sharing of resources and the cost of acquisition, GPUs (50x higher cost than CPUs) and high-memory systems are generally not available.
This complicates the deployment of models involving LLMs that require high GPU costs while still maintaining a healthy margin for sellers and the platform. Moreover, LLMs have large inference and training times [16] and have problems scaling on large datasets [17, 18] making them unsuitable for deployment in latency-sensitive applications.
Due to the model refresh requirements from Section I-A4, models with smaller training and setup times are absolutely necessary, and minimal tuning is crucial to decrease the engineering effort.
I-B Scope and Contributions
@@
In this work, we focus on the challenges mentioned in the previous section. We limit ourselves to retrieving keyphrases based on item’s title and belonging to the items’ categorical populace under budgetary constraints, especially those keyphrases that are actively and frequently searched by buyers. The extraction is done in an unsupervised setting where the keyphrases for each item are unknown during training. In fact, we restrict the curation of keyphrases (more details in III-B) to include only those that have a high search volume (number of searches made by buyers) based on buyer searches. XMC models suffer from the biases mentioned in Section I-A2 and as a result tend to recommend tail keyphrases (Section I-A1). Our distinction is that by using categorical population dynamics of keyphrases, we decouple the keyphrases from item click-based engagement. This allows us to retain the essential bias towards head keyphrases (attractive to advertisers) while getting rid of the negative bias (against non-popular items which are the main target of advertisement).
Our contribution to this work is summarized as follows:
@@
  • An innovative graph-based extraction algorithm for keyphrase recommendation that is transparent and easy to interpret.
@@
  • The design of the algorithm and the process of data collection have been specifically geared toward mitigating click-based biases while maintaining the advertiser friendly head keyphrase bias under budgetary constraints.
@@
  • Provide a new robust framework for the evaluation of incremental impact of recommendation models in terms of relevance and diversity metrics.
  • A low latency and sustainable model that runs without GPUs and scales for daily inference and training on billions of items.
II Related Work
Bipartite graphs are extensively utilized across various fields to model user search behaviors from logs, such as query-url [19, 20] and query-ad graphs [21]. Typically, techniques applied to these bipartite graphs calculate query similarities according to the items they are linked with. This similarity information is subsequently used to propose queries for new items. Simrank++ [22] enhances query similarity measures by reducing the iteration count required for convergence and by adjusting the similarity score with a multiplier that reflects the number of shared neighbors between queries. Nevertheless, in the worst-case scenario, these methods demand a pairwise comparison of all queries (i.e., quadratic complexity), which becomes impractical with a large volume of keyphrases. Additionally, ranking recommended queries linked to similar items based on their relevance presents another challenge.
SOTA models for extreme multi-label classification (XMC) [1, 2, 3, 4, 5] predominantly leverage deep neural networks (DNNs) and commonly employ one-vs-all (OVA) classifiers. Although certain models [1, 3] need label features, others such as AttentionXML [4], Renee [2], and DeepXML/Astec [5] do not. Among the DNN models assessed in [23], DeepXML/Astec [5] demonstrates scalability to large datasets and achieves a relatively short training duration compared to rival methods. However, [6] shows the non-viability of Astec and AttentionXML on large categories on eBay and the cost and scalability challenges of GPU inference associated with it.
Other efficient XMC tagging models include — fastText [24, 25] which is an effective CPU-based option for managing extensive workloads. fastText creates word embeddings using the CBOW model and employs a straightforward linear neural network model with hierarchical softmax to improve the efficiency of training and inference processes. A key reason for fastText’s effectiveness is its integration of subword information into the embeddings. The size of the model can be easily reduced to conserve storage using methods such as quantization [26], as well as pruning the vocabularies of important phrases and title words. Graphite [6] is another SOTA XMC model that uses bipartite graphs to map words/tokens to the data points and then map them to the labels associated with the data points. It is implemented for multi-core systems having infinitesimal training time and uses parallelization for real-time inferencing. It’s training and inference scales well for hundreds of millions of items and labels. SL-emb [27] uses embeddings of the item’s title to compare and find similar listings, and then recommend the related queries. SL-emb is a dense retrieval model whose inference is implemented in two stages, namely, embedding generation and ANN[28]. The SL-emb model does not need to be trained daily and is based on the hypothesis that semantically close items have similar keyphrases. The SL-emb model is trained on Recs data (similar item recommendation for a hero item), which is shown to mitigate some of the bias from Ads click data [29].
Rule-based heuristic models are also in production as simple models that can provide recommendations for existing popular items. Rules Engine (RE) is a simple technique that stores item-keyphrase associations based on their co-occurrences (associated with buyer activity) in the search logs during the last 30 days which is around 13% of all active items (item coverage). It recommends keyphrases only for items in which buyers have shown interest and not for any new items. This is a 100% recall model in which buyers’ interest is reflected back to them. SL-query is also a rule-based model based on the hypothesis that similar listings share similar queries. SL-query recommends the associated queries of listings that share a keyphrase with the seed item. Both SL models’ predictions are truncated from a higher number of predictions using a Jaccard coefficient [30] threshold to ensure relevance of the predictions. The RE and SL-query models have a low item-coverage (since items with query associations are quite sparse as described in Section I-A2) and don’t offer recommendations on cold items. The implementation of the RE and SL techniques also employs a few other details, which we cannot discuss due to proprietary constraints.
@@ II Related Work
Keyphrase generation via open-vocabulary models like GROOV [31], One2Seq [32, 33] and One2One [34, 35, 36] are susceptible to recommending keyphrases that are not part of the label space [7]. Another formulation for keyphrase recommendation is keyphrase extraction with methods such as keyBERT[37], which have conventionally treated keyphrase recommendation as a two-step problem: keyphrase generation and keyphrase ranking. The basic keyBERT module considers keyphrase generation as an n-gram-based permutation problem, i.e., it generates all possible n-grams for a given n-gram range. The keyphrase ranking module then orders them using an encoder-based ranker tuned on some domain-specific supervised signal. This simple generation framework presents two main issues: 1) the token space is limited by token adjacency and token presence in the item’s text; 2) the keyphrase should also be in the universe of queries that buyers are searching for; which this simple generation model does not ensure, as described in I-A4.222keyBERT can also use LLMs as generators, but their time complexity is substantial.
Table I shows how the various SOTA frameworks for keyphrase recommendation perform with the challenges mentioned previously. fastText, Graphite and SL-emb are all XMC tagging models deployed at eBay for seller-side keyphrase recommendations and are used in this study for comparison along with the rule-based models RE and SL-query.
Criteria Frameworks
XMC-tagging OOV GraphEx
Feasible daily batch
or real-time
prediction latency?
Click data debiasing ?
Susceptible to RE
De-duplication ?
100% targeting
in vocabulary
keyphrases ?
Focus on popular
keyphrases?
TABLE I: Table showing the comparative analysis of the capabilities of the various types of frameworks for keyphrase recommendation.
III GraphEx Model
@@ III-A Problem Formulation
For efficiently solving the recommendation problem, we use the formulation of a permutation problem that permutes the title strings to match a given set of keyphrases. Let us consider a title string with $l$ words in it. The goal is to generate permutations of different lengths from the $l$ words. Now, given a list of predefined keyphrases, the possible permutations of the title string are constrained to match the keyphrases. Therefore, each permutation can exactly match a keyphrase or be part of some keyphrases, but if a title token is not part of any keyphrase then it is ignored. Thus, it does not limit the permutations to token adjacency or token presence in the item’s text. A naive brute force method is to generate all possible permutations of the $l$ words which will take $O(l!)$ time. Each keyphrase can be validated using hashing and string comparisons (each word can be an integer) and thus can take overall $O(l\times l!)$ time. This is infeasible to perform in real-time with limited amount of resources.
III-B Dataset Curation
@@
We aggregate our keyphrase datasets and their category associations from the search logs generated during buyer sessions on eBay.com. The keyphrases that buyers input during the search sessions are curated based on certain criteria, which we discuss here. The categorization of items are in form of a tree. The root level category is called as meta category (top level category) which branches down to the Leaf Category. eBay’s search engine Cassini shows a sufficient number of items (Recall Count) for each input query in its search results. Cassini determines the leaf category of the keyphrase and it is the same as the top-ranked item’s leaf category (lowest-level product categorization). Note that the item-keyphrase click-based associations are not curated into our datasets.
We restrict the number of curated keyphrases by only considering those that are heavily searched by the buyers. The number of times a keyphrase is queried is termed as (Search Count). The search count is used to explicitly control the proportions of head keyphrases in our dataset. The absolute values of Recall Count and Search Count are not essential in fact, an ranking (integer value) also works. All the unique keyphrases are aggregated for each meta category and are grouped for each leaf category within the meta category. Each keyphrase is associated with a Search Count and a Recall Count. Note that a keyphrase can be duplicated across different Leaf Categories.
III-C Terms and Notations